# A lower bound for the Weisfeiler-Leman dimension of circulant graphs

### v1

@@ 6.2. Inner multipliers
where $\pi=\pi_{\mathfrak{s}}$. Put $Y=X^{{\mathfrak{m}}^{\prime}}$, where ${\mathfrak{m}}^{\prime}\in\operatorname{Aut}(G)$ is such that $({\mathfrak{m}}^{\prime})^{{\mathfrak{s}}}={\mathfrak{m}}_{\mathfrak{s}}$. By Theorem 5.1(1), we have $Y\in{\cal S}$. Furthermore, $\langle{X}\rangle=\langle{Y}\rangle$ and $\operatorname{rad}(X)=\operatorname{rad}(Y)$. Therefore, ${\mathfrak{s}}={\mathfrak{s}}_{X}={\mathfrak{s}}_{Y}$ and $Y_{\mathfrak{s}}={X_{\mathfrak{s}}}^{{\mathfrak{m}}_{\mathfrak{s}}}$. Thus,
$X^{f}=\pi^{-1}({X_{\mathfrak{s}}}^{{\mathfrak{m}}_{\mathfrak{s}}})=\pi^{-1}({Y_{\mathfrak{s}}})=Y\in{\cal S},$
and ${\cal A}^{f}={\cal A}$, as required. ∎
Each normalized isomorphism $f\in\operatorname{Iso}({\cal A})$ leaves every ${\cal A}$-subgroup fixed (because the S-ring ${\cal A}$ is circulant). This yields the restriction isomorphism $f^{\mathfrak{s}}\in\operatorname{Iso}({\cal A}_{\mathfrak{s}})$ for every ${\cal A}$-section ${\mathfrak{s}}$. Note that the collection ${\mathfrak{m}}_{f}=\{f^{\mathfrak{s}}:\ {\mathfrak{s}}\in{\mathfrak{S}}_{0}\}$ is not necessarily an inner ${\cal A}$-multiplier (because $f^{\mathfrak{s}}$ does not always belong to $\operatorname{Aut}({\mathfrak{s}})$). Put
$\operatorname{{\mathfrak{M}}}_{aut}({\cal A})=\{{\mathfrak{m}}_{f}\in\operatorname{{\mathfrak{M}}}({\cal A}):\ f\in\operatorname{Aut}({\cal A})\}.$
Assume that the group $G$ is of odd order. Let ${\mathfrak{s}}\in{\mathfrak{S}}_{0}$. Since the S-ring ${\cal A}$ is totally normal, the S-ring ${\cal A}_{\mathfrak{s}}$ is normal and by Lemma 5.2(2), the isomorphism $f^{\mathfrak{s}}$ belongs to $\operatorname{Aut}({\mathfrak{s}})$. Thus, ${\mathfrak{m}}_{f}\in\operatorname{{\mathfrak{M}}}({\cal A})$ and $\operatorname{{\mathfrak{M}}}_{aut}({\cal A})$ is a subgroup of $\operatorname{{\mathfrak{M}}}({\cal A})$.
@@ 6.3. Algebraic fusions
Let ${\cal A}$ be a coset S-ring, i.e., ${\cal A}={\cal A}_{0}$, and let ${\mathfrak{M}}$ be a group of inner ${\cal A}$-multipliers. This group acts on the basic sets of ${\cal A}$, namely, if $X\in{\cal S}$ and ${\mathfrak{m}}\in{\mathfrak{M}}$, then
@@
where ${\mathfrak{s}}={\mathfrak{s}}_{X}$. We define the algebraic fusion of ${\cal A}$ with respect to the group ${\mathfrak{M}}$ as an S-ring ${\cal A}^{\mathfrak{M}}$ over $G$ with the basic sets
(23) $X^{\mathfrak{M}}=\bigcup_{{\mathfrak{m}}\in{\mathfrak{M}}}X^{\mathfrak{m}},\qquad X\in{\cal S}.$
In the sense of [5, Subsection 2.3.3], the Cayley scheme ${\cal X}({\cal A}^{\mathfrak{M}})$ is the algebraic fusion of the Cayley scheme ${\cal X}({\cal A})$ with respect to the group of the algebraic automorphisms $\varphi_{{\mathfrak{c}}_{\mathfrak{m}}}$ corresponding to the outer ${\cal A}$-multiplier ${\mathfrak{c}}_{\mathfrak{m}}=\{\{{\mathfrak{m}}_{\mathfrak{s}}\}:\ {\mathfrak{s}}\in{\mathfrak{S}}_{0}\}$, ${\mathfrak{m}}\in{\mathfrak{M}}$. Since the S-ring ${\cal A}$ is a coset one, formula (14) implies that
(24) $({\cal A}^{\mathfrak{M}})_{\mathfrak{s}}=({\cal A}_{\mathfrak{s}})^{{\mathfrak{M}}_{\mathfrak{s}}}=({\mathbb{Z}}{\mathfrak{s}})^{{\mathfrak{M}}_{\mathfrak{s}}}=\operatorname{cyc}({\mathfrak{M}}_{\mathfrak{s}},{\mathfrak{s}}).$
@@
Obviously, ${\cal A}^{\mathfrak{M}}\leq{\cal A}$. Therefore each ${\cal A}^{\mathfrak{M}}$-subgroup is an ${\cal A}$-subgroup. Conversely, if $H$ is an ${\cal A}$-subgroup, then by Lemma 6.1, we have $H^{\mathfrak{m}}=H^{f_{\mathfrak{m}}}=H$ for all ${\mathfrak{m}}\in{\mathfrak{M}}$, and hence $H$ is an ${\cal A}^{\mathfrak{M}}$-subgroup. Thus the S-rings ${\cal A}$ and ${\cal A}^{\mathfrak{M}}$ have the same ${\cal A}$-subgroups and
(25) ${\mathfrak{S}}({\cal A})={\mathfrak{S}}({\cal A}^{\mathfrak{M}}).$
Moreover, it is not hard to see that the relations $\preceq$ and $\sim$ on the ${\cal A}$- and ${\cal A}^{\mathfrak{M}}$-sections are the same.
@@
In the above notation, assume that $G$ is of odd order and the S-ring $\operatorname{cyc}({\mathfrak{M}}_{\mathfrak{s}},{\mathfrak{s}})$ is normal for all ${\mathfrak{s}}\in{\mathfrak{S}}_{0}({\cal A})$. Then ${\mathfrak{M}}\leq{\mathfrak{M}}_{aut}({\cal A}^{\mathfrak{M}})$ and
(26) ${\mathfrak{S}}_{0}({\cal A}^{\mathfrak{M}})={\mathfrak{S}}_{0}({\cal A})\quad\text{and}\quad({\cal A}^{\mathfrak{M}})_{0}={\cal A}.$
@@
To prove the first inclusion, we need to verify that for each ${\mathfrak{m}}\in{\mathfrak{M}}$, there exists $f\in\operatorname{Aut}({\cal A}^{\mathfrak{M}})$ such that ${\mathfrak{m}}={\mathfrak{m}}_{f}$. To this end, denote by $f$ the isomorphism $f_{\mathfrak{m}}\in\operatorname{Iso}({\cal A})$ from Lemma 6.1 (note that in our case, ${\cal A}={\cal A}_{0}$). Then ${\mathfrak{m}}={\mathfrak{m}}_{f}$. Moreover, for all $X\in{\cal S}$, we have
@@
$=\bigcup_{{\mathfrak{n}}\in{\mathfrak{M}}}\pi_{\mathfrak{s}}^{-1}({X_{\mathfrak{s}}}^{({\mathfrak{n}}{\mathfrak{m}})_{\mathfrak{s}}})=\bigcup_{{\mathfrak{n}}^{\prime}\in{\mathfrak{M}}}\pi_{\mathfrak{s}}^{-1}({X_{\mathfrak{s}}}^{({\mathfrak{n}}^{\prime})_{\mathfrak{s}}})=\bigcup_{{\mathfrak{n}}^{\prime}\in{\mathfrak{M}}}X^{{\mathfrak{n}}^{\prime}}=X^{\mathfrak{M}},$
where ${\mathfrak{s}}={\mathfrak{s}}_{X}$, see formulas (22) and (23). Thus, $f\in\operatorname{Aut}({\cal A}^{\mathfrak{M}})$, as required.
First, we prove that ${\mathfrak{S}}_{0}({\cal A})\subseteq{\mathfrak{S}}_{0}({\cal A}^{\mathfrak{M}})$. Since the relations $\preceq$ and $\sim$ on the ${\cal A}$- and ${\cal A}^{\mathfrak{M}}$-sections are the same, it suffices to verify that every principal ${\cal A}$-section ${\mathfrak{s}}$ is a principal ${\cal A}^{\mathfrak{M}}$-section. Suppose that ${\mathfrak{s}}={\mathfrak{s}}_{X}$ for some $X\in{\cal S}$. Then obviously $\langle{X}\rangle=\langle{X^{\mathfrak{M}}}\rangle$ and $\operatorname{rad}(X)\leq\operatorname{rad}(X^{\mathfrak{M}})$. Suppose on the contrary that $\operatorname{rad}(X)\neq\operatorname{rad}(X^{\mathfrak{M}})$. Then the radical of the set $(X^{\mathfrak{M}})_{\mathfrak{s}}\leq{\mathfrak{s}}$ is nontrivial. Since the group ${\mathfrak{s}}$ is of odd order and $(X^{\mathfrak{M}})_{\mathfrak{s}}$ is a highest basic set of the S-ring $({\cal A}^{\mathfrak{M}})_{\mathfrak{s}}=\operatorname{cyc}({\mathfrak{M}}_{\mathfrak{s}},{\mathfrak{s}})$, this S-ring is not normal by [9, Theorem 6.1], a contradiction.
From formula (14), it follows that ${\mathfrak{S}}_{0}({\cal A}^{\mathfrak{M}})\subseteq{\mathfrak{S}}_{0}(({\cal A}^{\mathfrak{M}})_{0})$. On the other hand, ${\cal A}$ is a coset S-ring containing ${\cal A}^{\mathfrak{M}}$. Therefore, $({\cal A}^{\mathfrak{M}})_{0}\leq{\cal A}$. It follows that ${\mathfrak{S}}_{0}(({\cal A}^{\mathfrak{M}})_{0})\subseteq{\mathfrak{S}}_{0}({\cal A})$. Consequently, ${\mathfrak{S}}_{0}({\cal A}^{\mathfrak{M}})\subseteq{\mathfrak{S}}_{0}({\cal A})$. Together with the inclusion in the above paragraph, this proves the first equality in (26). Furthermore, the above argument shows that ${\mathfrak{S}}_{0}(({\cal A}^{\mathfrak{M}})_{0})={\mathfrak{S}}_{0}({\cal A})$. Consequently, the coset S-rings $({\cal A}^{\mathfrak{M}})_{0}$ and ${\cal A}$ have the same principal sections. Thus they are equal which proves the second equality in (26). ∎
@@ 7. Local system of multipliers
Let ${\cal A}$ be a circulant S-ring, ${\mathfrak{c}}$ an outer ${\cal A}$-multiplier, and $k\geq 2$ an integer. Denote by ${\mathfrak{S}}_{k}={\mathfrak{S}}_{k}({\cal A})$ the set of all subsets consisting of no more than $k$ principal ${\cal A}$-sections. Assume that for each ${\mathfrak{S}}\in{\mathfrak{S}}_{k}$, we are given an inner ${\cal A}$-multiplier ${\mathfrak{m}}({\mathfrak{S}})$. A family
(27) $\{{\mathfrak{m}}({\mathfrak{S}})\in{\mathfrak{M}}({\cal A}):\ {\mathfrak{S}}\in{\mathfrak{S}}_{k}\}$
is called a $k$-local system of inner ${\cal A}$-multipliers with respect to ${\mathfrak{c}}$ if the following two conditions are satisfied for all subsets ${\mathfrak{S}},{\mathfrak{S}}^{\prime}\in{\mathfrak{S}}_{k}$ and some inner ${\cal A}$-multipliers ${\mathfrak{m}}({\mathfrak{S}},{\mathfrak{S}}^{\prime})\in\operatorname{{\mathfrak{M}}}_{aut}({\cal A})$:
@@
Let ${\cal A}$ be a totally normal circulant S-ring and $k\geq 2$ an integer. Assume that
@@
    the scheme ${\cal X}({\cal A}_{0})$ is binary,
@@
    there is a $k$-local system of inner ${\cal A}$-multipliers with respect to an outer multiplier ${\mathfrak{c}}\in{\mathfrak{C}}({\cal A})$.
Then for each $2\leq m<\sqrt{k}$, the scheme ${\cal X}({\cal A})$ is $\operatorname{WL}_{m}$-equivalent to itself with respect to the algebraic isomorphism $\varphi_{\mathfrak{c}}$.
We prove Theorem 7.1 at the end of the section. In what follows, $G$ is a cyclic group underlying the S-ring ${\cal A}$, ${\cal X}={\cal X}({\cal A})$ is the Cayley scheme associated with ${\cal A}$, and $\varphi=\varphi_{\mathfrak{c}}$. Fix an integer $\ell$ such that
@@
and fix a $k$-local system (27) of inner ${\cal A}$-multipliers with respect to the outer multiplier ${\mathfrak{c}}$.
Denote by ${\cal R}_{\ell}={\cal R}_{\ell}({\cal X})$ the partition of the Cartesian power $G^{\ell}$ into the classes of the equivalence relation $\sim_{\ell}$ defined as follows: $x\sim_{\ell}y$ if and only if $R_{\ell}(x)=R_{\ell}(y)$, where $R_{\ell}(\cdot)=R_{{\cal X},\ell}(\cdot)$ is the $\ell\times\ell$-array defined in formula (7). In particular, each class ${\mathscr{X}}\in{\cal R}_{\ell}$ is a maximal subset of $G^{\ell}$ such that the array $R_{\ell}(x)$ does not depend on the choice of the $\ell$-tuple $x\in{\mathscr{X}}$. In accordance with [6, Subsection 3.1], we have
(30) $\operatorname{WL}_{\ell}({\cal X})\geq{\cal R}_{\ell}({\cal X}),$
i.e., each class of the partition ${\cal R}_{\ell}({\cal X})$ is a union of some classes of the partition $\operatorname{WL}_{\ell}({\cal X})$. Furthermore, the algebraic isomorphism $\varphi$ induces a permutation of ${\cal R}_{\ell}({\cal X})$, that takes ${\mathscr{X}}$ to ${\mathscr{X}}^{\varphi}=\{y\in G^{\ell}:\ R(y)=R(x)^{\varphi}\}$, where $R(x)^{\varphi}=(r(x_{i},x_{j})^{\varphi})$ and $x\in{\mathscr{X}}$. In particular,
(31) ${\cal R}_{\ell}({\cal X})^{\hat{\varphi}}={\cal R}_{\ell}({\cal X}).$
Let ${\mathscr{X}}$ be a class of the partition ${\cal R}_{\ell}$, and let $1\leq i,j\leq\ell$. The basis relation $s=r(x_{i},x_{j})$ of the scheme ${\cal X}$ does not depend on the $\ell$-tuple $x\in{\mathscr{X}}$. Denote by $X_{ij}$ the basic set of the S-ring ${\cal A}$, corresponding to $s$. Clearly, $x_{j}{x_{i}}^{-1}\in X_{ij}$ for all $x\in{\mathscr{X}}$. The set
[…]
consists of at most $\ell^{2}\leq k$ principal ${\cal A}$-sections ${\mathfrak{s}}_{X_{ij}}$ and hence ${\mathfrak{S}}({\mathscr{X}})\in{\mathfrak{S}}_{k}$. Denote by ${\mathfrak{m}}({\mathscr{X}})$ the inner ${\cal A}$-multiplier ${\mathfrak{m}}({\mathfrak{S}})$ belonging to the family (27) for ${\mathfrak{S}}={\mathfrak{S}}({\mathscr{X}})$, and by $f_{\mathscr{X}}$ the isomorphism $f_{{\mathfrak{m}}({\mathscr{X}})}\in\operatorname{Iso}_{\cal A}({\cal A}_{0})$ in Lemma 6.1 for ${\mathfrak{m}}={\mathfrak{m}}({\mathscr{X}})$.
@@
[…]
(33) ${\cal R}_{\ell}({\cal X})^{f}={\cal R}_{\ell}({\cal X})\quad\text{and}\quad\operatorname{WL}_{\ell}({\cal X})^{f}=\operatorname{WL}_{\ell}({\cal X}).$

### v2

@@ 6.2. Inner multipliers
where $\pi=\pi_{\mathfrak{s}}$. Put $Y=X^{{\mathfrak{m}}^{\prime}}$, where ${\mathfrak{m}}^{\prime}\in\operatorname{Aut}(G)$ is such that $({\mathfrak{m}}^{\prime})^{{\mathfrak{s}}}={\mathfrak{m}}_{\mathfrak{s}}$. By Theorem 5.1(1), we have $Y\in{\mathcal{S}}$. Furthermore, $\langle{X}\rangle=\langle{Y}\rangle$ and $\operatorname{rad}(X)=\operatorname{rad}(Y)$. Therefore, ${\mathfrak{s}}={\mathfrak{s}}_{X}={\mathfrak{s}}_{Y}$ and $Y_{\mathfrak{s}}={X_{\mathfrak{s}}}^{{\mathfrak{m}}_{\mathfrak{s}}}$. Thus,
$X^{f}=\pi^{-1}({X_{\mathfrak{s}}}^{{\mathfrak{m}}_{\mathfrak{s}}})=\pi^{-1}({Y_{\mathfrak{s}}})=Y\in{\mathcal{S}},$
and ${\mathcal{A}}^{f}={\mathcal{A}}$, as required. ∎
Each normalized isomorphism $f\in\operatorname{Iso}({\mathcal{A}})$ leaves every $\mathcal{A}$}-subgroup fixed (because the S-ring $\mathcal{A}$} is circulant). This yields the restriction isomorphism $f^{\mathfrak{s}}\in\operatorname{Iso}({\mathcal{A}}_{\mathfrak{s}})$ for every $\mathcal{A}$}-section $\mathfrak{s}$}. Note that the collection ${\mathfrak{m}}_{f}=\{f^{\mathfrak{s}}:\ {\mathfrak{s}}\in{\mathfrak{S}}_{0}\}$ is not necessarily an inner $\mathcal{A}$}-multiplier (because $f^{\mathfrak{s}}$ does not always belong to $\operatorname{Aut}({\mathfrak{s}})$). Put
$\operatorname{{\mathfrak{M}}}_{aut}({\mathcal{A}})=\{{\mathfrak{m}}_{f}\in\operatorname{{\mathfrak{M}}}({\mathcal{A}}):\ f\in\operatorname{Aut}({\mathcal{A}})\}.$
Assume that the group $G$ is of odd order. Let ${\mathfrak{s}}\in{\mathfrak{S}}_{0}$. Since the S-ring $\mathcal{A}$} is totally normal, the S-ring ${\mathcal{A}}_{\mathfrak{s}}$ is normal and by Lemma 5.2(2), the isomorphism $f^{\mathfrak{s}}$ belongs to $\operatorname{Aut}({\mathfrak{s}})$. Thus, ${\mathfrak{m}}_{f}\in\operatorname{{\mathfrak{M}}}({\mathcal{A}})$ and $\operatorname{{\mathfrak{M}}}_{aut}({\mathcal{A}})$ is a subgroup of $\operatorname{{\mathfrak{M}}}({\mathcal{A}})$.
@@ 6.3. Algebraic fusions
Let $\mathcal{A}$} be a coset S-ring, i.e., ${\mathcal{A}}={\mathcal{A}}_{0}$, and let $\mathfrak{M}$} be a group of inner $\mathcal{A}$}-multipliers. This group acts on the basic sets of $\mathcal{A}$}, namely, if $X\in{\mathcal{S}}$ and ${\mathfrak{m}}\in{\mathfrak{M}}$, then
@@
where ${\mathfrak{s}}={\mathfrak{s}}_{X}$. We define the algebraic fusion of $\mathcal{A}$} with respect to the group $\mathfrak{M}$} as an S-ring ${\mathcal{A}}^{\mathfrak{M}}$ over $G$ with the basic sets
(23) $X^{\mathfrak{M}}=\bigcup_{{\mathfrak{m}}\in{\mathfrak{M}}}X^{\mathfrak{m}},\qquad X\in{\mathcal{S}}.$
In the sense of [5, Subsection 2.3.3], the Cayley scheme ${\mathcal{X}}({\mathcal{A}}^{\mathfrak{M}})$ is the algebraic fusion of the Cayley scheme ${\mathcal{X}}({\mathcal{A}})$ with respect to the group of the algebraic automorphisms $\varphi_{{\mathfrak{c}}_{\mathfrak{m}}}$ corresponding to the outer $\mathcal{A}$}-multiplier ${\mathfrak{c}}_{\mathfrak{m}}=\{\{{\mathfrak{m}}_{\mathfrak{s}}\}:\ {\mathfrak{s}}\in{\mathfrak{S}}_{0}\}$, ${\mathfrak{m}}\in{\mathfrak{M}}$. Since the S-ring $\mathcal{A}$} is a coset one, formula (14) implies that
(24) $({\mathcal{A}}^{\mathfrak{M}})_{\mathfrak{s}}=({\mathcal{A}}_{\mathfrak{s}})^{{\mathfrak{M}}_{\mathfrak{s}}}=({\mathbb{Z}}{\mathfrak{s}})^{{\mathfrak{M}}_{\mathfrak{s}}}=\operatorname{cyc}({\mathfrak{M}}_{\mathfrak{s}},{\mathfrak{s}}).$
@@
Obviously, ${\mathcal{A}}^{\mathfrak{M}}\leq{\mathcal{A}}$. Therefore each ${\mathcal{A}}^{\mathfrak{M}}$-subgroup is an $\mathcal{A}$}-subgroup. Conversely, if $H$ is an $\mathcal{A}$}-subgroup, then by Lemma 6.1, we have $H^{\mathfrak{m}}=H^{f_{\mathfrak{m}}}=H$ for all ${\mathfrak{m}}\in{\mathfrak{M}}$, and hence $H$ is an ${\mathcal{A}}^{\mathfrak{M}}$-subgroup. Thus the S-rings $\mathcal{A}$} and ${\mathcal{A}}^{\mathfrak{M}}$ have the same $\mathcal{A}$}-subgroups and
(25) ${\mathfrak{S}}({\mathcal{A}})={\mathfrak{S}}({\mathcal{A}}^{\mathfrak{M}}).$
Moreover, it is not hard to see that the relations $\preceq$ and $\sim$ on the $\mathcal{A}$}- and ${\mathcal{A}}^{\mathfrak{M}}$-sections are the same.
@@
In the above notation, assume that $G$ is of odd order and the S-ring $\operatorname{cyc}({\mathfrak{M}}_{\mathfrak{s}},{\mathfrak{s}})$ is normal for all ${\mathfrak{s}}\in{\mathfrak{S}}_{0}({\mathcal{A}})$. Then ${\mathfrak{M}}\leq{\mathfrak{M}}_{aut}({\mathcal{A}}^{\mathfrak{M}})$ and
(26) ${\mathfrak{S}}_{0}({\mathcal{A}}^{\mathfrak{M}})={\mathfrak{S}}_{0}({\mathcal{A}})\quad\text{and}\quad({\mathcal{A}}^{\mathfrak{M}})_{0}={\mathcal{A}}.$
@@
To prove the first inclusion, we need to verify that for each ${\mathfrak{m}}\in{\mathfrak{M}}$, there exists $f\in\operatorname{Aut}({\mathcal{A}}^{\mathfrak{M}})$ such that ${\mathfrak{m}}={\mathfrak{m}}_{f}$. To this end, denote by $f$ the isomorphism $f_{\mathfrak{m}}\in\operatorname{Iso}({\mathcal{A}})$ from Lemma 6.1 (note that in our case, ${\mathcal{A}}={\mathcal{A}}_{0}$). Then ${\mathfrak{m}}={\mathfrak{m}}_{f}$. Moreover, for all $X\in{\mathcal{S}}$, we have
@@
$\bigcup_{{\mathfrak{n}}\in{\mathfrak{M}}}\pi_{\mathfrak{s}}^{-1}({X_{\mathfrak{s}}}^{({\mathfrak{n}}{\mathfrak{m}})_{\mathfrak{s}}})=\bigcup_{{\mathfrak{n}}^{\prime}\in{\mathfrak{M}}}\pi_{\mathfrak{s}}^{-1}({X_{\mathfrak{s}}}^{({\mathfrak{n}}^{\prime})_{\mathfrak{s}}})=\bigcup_{{\mathfrak{n}}^{\prime}\in{\mathfrak{M}}}X^{{\mathfrak{n}}^{\prime}}=X^{\mathfrak{M}},$
where ${\mathfrak{s}}={\mathfrak{s}}_{X}$, see formulas (22) and (23). Thus, $f\in\operatorname{Aut}({\mathcal{A}}^{\mathfrak{M}})$, as required.
First, we prove that ${\mathfrak{S}}_{0}({\mathcal{A}})\subseteq{\mathfrak{S}}_{0}({\mathcal{A}}^{\mathfrak{M}})$. Since the relations $\preceq$ and $\sim$ on the $\mathcal{A}$}- and ${\mathcal{A}}^{\mathfrak{M}}$-sections are the same, it suffices to verify that every principal $\mathcal{A}$}-section $\mathfrak{s}$} is a principal ${\mathcal{A}}^{\mathfrak{M}}$-section. Suppose that ${\mathfrak{s}}={\mathfrak{s}}_{X}$ for some $X\in{\mathcal{S}}$. Then obviously $\langle{X}\rangle=\langle{X^{\mathfrak{M}}}\rangle$ and $\operatorname{rad}(X)\leq\operatorname{rad}(X^{\mathfrak{M}})$. Suppose on the contrary that $\operatorname{rad}(X)\neq\operatorname{rad}(X^{\mathfrak{M}})$. Then the radical of the set $(X^{\mathfrak{M}})_{\mathfrak{s}}\leq{\mathfrak{s}}$ is nontrivial. Since the group $\mathfrak{s}$} is of odd order and $(X^{\mathfrak{M}})_{\mathfrak{s}}$ is a highest basic set of the S-ring $({\mathcal{A}}^{\mathfrak{M}})_{\mathfrak{s}}=\operatorname{cyc}({\mathfrak{M}}_{\mathfrak{s}},{\mathfrak{s}})$, this S-ring is not normal by [9, Theorem 6.1], a contradiction.
From formula (14), it follows that ${\mathfrak{S}}_{0}({\mathcal{A}}^{\mathfrak{M}})\subseteq{\mathfrak{S}}_{0}(({\mathcal{A}}^{\mathfrak{M}})_{0})$. On the other hand, $\mathcal{A}$} is a coset S-ring containing ${\mathcal{A}}^{\mathfrak{M}}$. Therefore, $({\mathcal{A}}^{\mathfrak{M}})_{0}\leq{\mathcal{A}}$. It follows that ${\mathfrak{S}}_{0}(({\mathcal{A}}^{\mathfrak{M}})_{0})\subseteq{\mathfrak{S}}_{0}({\mathcal{A}})$. Consequently, ${\mathfrak{S}}_{0}({\mathcal{A}}^{\mathfrak{M}})\subseteq{\mathfrak{S}}_{0}({\mathcal{A}})$. Together with the inclusion in the above paragraph, this proves the first equality in (26). Furthermore, the above argument shows that ${\mathfrak{S}}_{0}(({\mathcal{A}}^{\mathfrak{M}})_{0})={\mathfrak{S}}_{0}({\mathcal{A}})$. Consequently, the coset S-rings $({\mathcal{A}}^{\mathfrak{M}})_{0}$ and $\mathcal{A}$} have the same principal sections. Thus they are equal which proves the second equality in (26). ∎
@@ 7. Local system of multipliers
Let $\mathcal{A}$} be a circulant S-ring, $\mathfrak{c}$} an outer $\mathcal{A}$}-multiplier, and $k\geq 2$ an integer. Denote by ${\mathfrak{S}}_{k}={\mathfrak{S}}_{k}({\mathcal{A}})$ the set of all subsets consisting of no more than $k$ principal $\mathcal{A}$}-sections. Assume that for each ${\mathfrak{S}}\in{\mathfrak{S}}_{k}$, we are given an inner $\mathcal{A}$}-multiplier ${\mathfrak{m}}({\mathfrak{S}})$. A family
(27) $\{{\mathfrak{m}}({\mathfrak{S}})\in{\mathfrak{M}}({\mathcal{A}}):\ {\mathfrak{S}}\in{\mathfrak{S}}_{k}\}$
is called a $k$-local system of inner $\mathcal{A}$}-multipliers with respect to $\mathfrak{c}$} if the following two conditions are satisfied for all subsets ${\mathfrak{S}},{\mathfrak{S}}^{\prime}\in{\mathfrak{S}}_{k}$ and some inner $\mathcal{A}$}-multipliers ${\mathfrak{m}}({\mathfrak{S}},{\mathfrak{S}}^{\prime})\in\operatorname{{\mathfrak{M}}}_{aut}({\mathcal{A}})$:
@@
Let $\mathcal{A}$} be a totally normal circulant S-ring and $k\geq 2$ an integer. Assume that
@@
    the scheme ${\mathcal{X}}({\mathcal{A}}_{0})$ is binary,
@@
    there is a $k$-local system of inner $\mathcal{A}$}-multipliers with respect to an outer multiplier ${\mathfrak{c}}\in{\mathfrak{C}}({\mathcal{A}})$.
Then for each $2\leq m<\sqrt{k}$, the scheme ${\mathcal{X}}({\mathcal{A}})$ is $\operatorname{WL}_{m}$-equivalent to itself with respect to the algebraic isomorphism $\varphi_{\mathfrak{c}}$.
We prove Theorem 7.1 at the end of the section. In what follows, $G$ is a cyclic group underlying the S-ring $\mathcal{A}$}, ${\mathcal{X}}={\mathcal{X}}({\mathcal{A}})$ is the Cayley scheme associated with $\mathcal{A}$}, and $\varphi=\varphi_{\mathfrak{c}}$. Fix an integer $\ell$ such that
@@
and fix a $k$-local system (27) of inner $\mathcal{A}$}-multipliers with respect to the outer multiplier $\mathfrak{c}$}.
Denote by ${\mathcal{R}}_{\ell}={\mathcal{R}}_{\ell}({\mathcal{X}})$ the partition of the Cartesian power $G^{\ell}$ into the classes of the equivalence relation $\sim_{\ell}$ defined as follows: $x\sim_{\ell}y$ if and only if $R_{\ell}(x)=R_{\ell}(y)$, where $R_{\ell}(\cdot)=R_{{\mathcal{X}},\ell}(\cdot)$ is the $\ell\times\ell$-array defined in formula (7). In particular, each class $\mathscr{X}$$\in{\mathcal{R}}_{\ell}$ is a maximal subset of $G^{\ell}$ such that the array $R_{\ell}(x)$ does not depend on the choice of the $\ell$-tuple $x\in{\mathscr{X}}$. In accordance with [6, Subsection 3.1], we have
(30) $\operatorname{WL}_{\ell}({\mathcal{X}})\geq{\mathcal{R}}_{\ell}({\mathcal{X}}),$
i.e., each class of the partition ${\mathcal{R}}_{\ell}({\mathcal{X}})$ is a union of some classes of the partition $\operatorname{WL}_{\ell}({\mathcal{X}})$. Furthermore, the algebraic isomorphism $\varphi$ induces a permutation of ${\mathcal{R}}_{\ell}({\mathcal{X}})$, that takes $\mathscr{X}$} to ${\mathscr{X}}^{\varphi}=\{y\in G^{\ell}:\ R(y)=R(x)^{\varphi}\}$, where $R(x)^{\varphi}=(r(x_{i},x_{j})^{\varphi})$ and $x\in{\mathscr{X}}$. In particular,
(31) ${\mathcal{R}}_{\ell}({\mathcal{X}})^{\hat{\varphi}}={\mathcal{R}}_{\ell}({\mathcal{X}}).$
Let $\mathscr{X}$} be a class of the partition $\mathcal{R}$$_{\ell}$, and let $1\leq i,j\leq\ell$. The basis relation $s=r(x_{i},x_{j})$ of the scheme $\mathcal{X}$} does not depend on the $\ell$-tuple $x\in{\mathscr{X}}$. Denote by $X_{ij}$ the basic set of the S-ring $\mathcal{A}$}, corresponding to $s$. Clearly, $x_{j}{x_{i}}^{-1}\in X_{ij}$ for all $x\in{\mathscr{X}}$. The set
(32) ${\mathfrak{S}}({\mathscr{X}})=\{{\mathfrak{s}}_{X_{ij}}:\ 1\leq i,j\leq\ell\}$
consists of at most $\ell^{2}\leq k$ principal $\mathcal{A}$}-sections ${\mathfrak{s}}_{X_{ij}}$ and hence ${\mathfrak{S}}({\mathscr{X}})\in{\mathfrak{S}}_{k}$. Denote by ${\mathfrak{m}}({\mathscr{X}})$ the inner $\mathcal{A}$}-multiplier ${\mathfrak{m}}({\mathfrak{S}})$ belonging to the family (27) for ${\mathfrak{S}}={\mathfrak{S}}({\mathscr{X}})$, and by $f_{\mathscr{X}}$ the isomorphism $f_{{\mathfrak{m}}({\mathscr{X}})}\in\operatorname{Iso}_{\mathcal{A}}({\mathcal{A}}_{0})$ in Lemma 6.1 for ${\mathfrak{m}}={\mathfrak{m}}({\mathscr{X}})$.
@@
is a bijection, where ${\mathscr{X}}\in{\mathcal{R}}_{\ell}({\mathcal{X}})$ is the class containing the $\ell$-tuple $x$ and $x^{f_{\mathscr{X}}}$ denotes the $\ell$-tuple $(x_{1}^{f_{\mathscr{X}}},\ldots,x_{\ell}^{f_{\mathscr{X}}})$. Moreover,
(33) ${\mathcal{R}}_{\ell}({\mathcal{X}})^{f}={\mathcal{R}}_{\ell}({\mathcal{X}})\quad\text{and}\quad\operatorname{WL}_{\ell}({\mathcal{X}})^{f}=\operatorname{WL}_{\ell}({\mathcal{X}}).$